OpenStudy (anonymous):

Find GCD(-27,-45). I think that the answer is 9, but the back of the book says 3. Which do you think is correct and why?

OpenStudy (amistre64):

what are their prime factors?

OpenStudy (amistre64):

i dont think ive got this scripted for negative numbers ... so it might be best to enter them in absolute form

OpenStudy (anonymous):

Well I believe that with negative numbers, it is still the highest positive factor (which would be 9). I'm just not completely sure.

OpenStudy (amistre64):

you are correct, i just didnt have my scripting for the program to account for negative entries :)

OpenStudy (amistre64):

if -9 is a factor, then so is +9, and +9 is greater then -9

OpenStudy (skullpatrol):

Do you know how to find the prime factorization of a number?

OpenStudy (amistre64):

i find prime factors like a type ... hunt and peck :)

OpenStudy (skullpatrol):

@Skip2mylou426 ?

OpenStudy (anonymous):

I thought so until I hit number theory haha

OpenStudy (skullpatrol):

Divide by the prime numbers in order as many times as possible before going on to the next prime.

OpenStudy (anonymous):

I don't think it's asking for primes, I'm pretty sure it's just asking for the GCD

OpenStudy (skullpatrol):

Oops I meant the GCD is the GCF, right?

OpenStudy (amistre64):

if d|a and d|b then d is a common divisor the largest common divisor is the GCD (greatest common divisor)

OpenStudy (anonymous):

That's why I think it's 9, but I can't figure out why the back of the book says 3. That makes no sense to me.

OpenStudy (amistre64):

3 | -27 and -45 9 | -27 and -45 since 9 > 3, 9 is the gcd

OpenStudy (amistre64):

make sure your looking at the correct answer solution :)

OpenStudy (amistre64):

45 = 27(1) + 18 27 = 18(1) + 9 <-- gcd is 9 18 = 9(2) + 0

OpenStudy (skullpatrol):

Maybe it's a typo that should have been 3^2?

OpenStudy (phi):

if you haven't noticed, you will, that math books often have typos proof readers have a hard time picking up mistakes. You can use wolfram as a check on some of these questions http://www.wolframalpha.com/input/?i=gcd%28-27%2C-45%29

OpenStudy (skullpatrol):

Yes, even the best ones aren't typo-free.

OpenStudy (anonymous):

"The positive divisors of 27 are 1, 3, 9, and 27. The positive divisors of 45 are 1, 3, 5, 9, 15, and 45. It follows that the greatest positive divisor of 27 and 45 is 3."

OpenStudy (anonymous):

I was thinking that it might be a typo, but I'm looking at their lists and it's just not making sense. I'm gonna stick with 9. Thanks for your input everyone!

OpenStudy (amistre64):

good luck ;)